OpenStudy (anonymous):

sin(3x)[sin(2x)cos(x)+cos(2x)sin(x)]=3/4 in the interval [0,2pi)

OpenStudy (anonymous):

You must know that: \[\sin(A) \cos(B) + \cos(A) \sin(B) = \sin(A+B)\]

OpenStudy (anonymous):

Plug \(A = 2x\) and \(B = x\) in that formula.. :) And tell me what did you get?